A cylinder with 25.0 L chlorine gas initially had a pressure of 625 mmHg at 125 degree celsius. Changes in temperature caused volume of the gas to decrease 15,000 mL with new pressure of 1.5 atm. What was the change in temperature in degree celsius, that caused this change?
Answer #1

ch gas pressure = P, Volume = V, Temperature = 7 Initially Finally V = 25L V = 15,000ml = 15L P, = 6a5mm Hg T = 125°C =(125+2Change in temperature (Toz – T) = AT AT = 162-57°C - 195°C AT = 37.57% AT=37. 60